# Langevin equations and a geometric integration scheme for the overdamped limit of homogeneous rotational Brownian motion

### v2

@@ V Extension to axisymmetric particles
with constants $D_{R}^{\parallel,\perp}=k_{\text{B}}T/\zeta_{R}^{\parallel,\perp}$. A derivation and further explanations of Eqs. 50a and 50b can be found in Appendix A.
VI Summary and conclusions
@@
Brownian motion on the unit sphere is a central ingredient to models of axisymmetric colloidal particles, in which it describes the stochastic evolution of the symmetry axis; examples are oblate and prolate ellipsoids, but also active Brownian particles, e.g., Janus spheres, where such an axis is given by the propulsion direction. Here, we have constructed the corresponding overdamped stochastic process as a sequence of infinitesimal random rotations [Eq. 10]. This construction is known as McKean–Gangolli injection [83], which transforms a stochastic process on the Lie algebra — here, $\mathfrak{so}(3)$ — into the generated Lie group — here, the rotation group $SO(3)$. Within Itō calculus, it is then straightforward to derive a stochastic differential (i.e., Langevin) equation of the process [Eq. 9]. Transforming the equation to its equivalent form in Stratonovich calculus [Eq. 7] yields the same equation as is obtained heuristically from the underdamped Langevin equation of rotational Brownian motion [Eqs. 2 and 4]. In particular, the latter form is free from the apparent, noise-induced drift term present in the Itō form. We have shown that, indeed, this additional term does not generate a drift (bias) of the rotational motion; but rather, it is needed to preserve the normalisation of the orientation vector, $|\boldsymbol{u}(t)|=1$, under the stochastic dynamics.
@@
In generic numerical integration schemes of the overdamped Langevin equation (in either form), such as the Euler–Maruyama and Milstein schemes, the normalisation constraint is satisfied only asymptotically for vanishing time step, ${\Delta t\to 0}$, depending on their strong order of convergence. As an alternative, we have proposed a geometric integration scheme [Eq. 24], which is based on finite random rotations and which satisfies the constraint exactly for any time step $\Delta t$. Combining this approach with Gaussian distributed random rotation angles yields an immediate and simple implementation of this geometric integrator (Table 1), which we have shown to converge weakly at order 1, i.e., $O(\Delta t)$, and which can also be combined with a deterministic bias, e.g., due to an external torque or to model circle swimmers. Moreover, we have outlined an improved algorithm that exactly generates the (free) propagator of rotational Brownian motion for an arbitrary time step $\Delta t$ if an advanced sampling of the rotation angle is used [see Eq. 37]. Finally, we have demonstrated numerically and by deriving the corresponding propagators that the proposed Gaussian geometric scheme converges more rapidly than a widely used ad hoc scheme [Eqs. (38)]; the smaller prefactor of the discretisation error allows for an integration time step that is an order of magnitude larger (Fig. 1).
@@ Appendix A Statistics of the projected angular velocity
For a derivation of the first two moments of the projected angular velocity $\omega_{\perp}(t)$, given in Eqs. 50a and 50b, we switch to the integrated noise process,
@@
which in differential form reads $\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=\boldsymbol{\omega}_{\perp}(t)\,\mathrm{d}t$. It is a proper Itō processes and, using Eqs. 3 and 49, it obeys the following (Itō) SDE in standard form [51, 52]:
$\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\,\mathrm{d}t+(2D_{R}^{\perp})^{1/2}\mathcal{Q}_{\boldsymbol{u}}(t)\,\mathrm{d}\boldsymbol{W}(t)\,,$ (53)
which is driven by a standard three-dimensional Wiener process $\boldsymbol{W}(t)$ normalised such that $\langle|\boldsymbol{W}(t)|^{2}\rangle=3t$.
@@
$=E\biggl{[}\int_{t}^{t+\Delta t}(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})\boldsymbol{T}_{\text{ext}}(t^{\prime})\,\mathrm{d}t^{\prime}\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$\simeq(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\Delta t\,;\quad\Delta t\to 0.$ (54)
@@
Switching to the derivative of $\boldsymbol{\Omega}_{\perp}(t)$ in Eq. 54 yields Eq. 50a.
For the covariance function of $\boldsymbol{\Omega}_{\perp}(t)$, we compensate the drift and consider the rescaled process $\mathrm{d}\boldsymbol{W}^{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}}(t)\,\mathrm{d}\boldsymbol{W}(t)$ such that $\int_{0}^{t}\delta\boldsymbol{\omega}_{\perp}(t^{\prime})\,\mathrm{d}t^{\prime}=(2D_{R}^{\perp})^{1/2}\boldsymbol{W}^{\perp}(t)$. Non-overlapping increments of $\boldsymbol{W}^{\perp}(t)$ are uncorrelated, which is seen by choosing $\Delta t^{\prime},\Delta t>0$ and $t^{\prime}>t+\Delta t$ and performing basic manipulations of conditional expectation:
@@
$=0\,.$ (56)
Choosing the same time interval for both increments, it holds:
@@
$\simeq[\mathcal{Q}_{\boldsymbol{u}}(t)]_{ij}\Delta t\,;\quad\Delta t\to 0\,;$ (57)

### v3

@@ V Extension to axisymmetric particles
with constants $D_{R}^{\parallel,\perp}=k_{\text{B}}T/\zeta_{R}^{\parallel,\perp}$. A derivation and further explanations of Eqs.˜51a and 51b can be found in Appendix˜A.
VI Fokker–Planck equation
Before closing, we discuss the Fokker–Planck equation, which determines the propagator of the process and which is complementary to the Langevin description of the stochastic evolution of the trajectories [Eqs.˜7 and 9]. The (two-time) propagator $p(\boldsymbol{u},t|\boldsymbol{u}_{0},t_{0})$ is the probability density of the orientation $\boldsymbol{u}(t)$ at time $t$ given its value $\boldsymbol{u}_{0}=\boldsymbol{u}(t_{0})$ at an earlier time $t_{0}<t$. Due to the Markov property of diffusion, $p(\boldsymbol{u},t|\boldsymbol{u}_{0},t_{0})$ specifies the complete statistics of the dynamics (which is also referred to as the “probability law” of the process). The Fokker–Planck equation is a partial differential equation (PDE) with an elliptic differential operator $\mathcal{L}_{\boldsymbol{u}}(t)$ acting on the dependence on $\boldsymbol{u}$:
$\partial_{t}p(\boldsymbol{u},t|\boldsymbol{u}_{0},t_{0})=\mathcal{L}_{\boldsymbol{u}}(t)\,p(\boldsymbol{u},t|\boldsymbol{u}_{0},t_{0})$ (53)
for $t\geqslant t_{0}$ and subject to the initial condition $p(\boldsymbol{u},t_{0}|\boldsymbol{u}_{0},t_{0})=\delta(\boldsymbol{u}-\boldsymbol{u}_{0})$. For a time-homogeneous process, $\mathcal{L}_{\boldsymbol{u}}$ is constant and the propagator depends only on $t-t_{0}$, so we can set $t_{0}=0$ and omit this dependence. For free diffusion on a manifold, $\mathcal{L}_{\boldsymbol{u}}$ is the Laplace–Beltrami operator [52, 80], which in case of the unit sphere is simply the angular part of the three-dimensional Laplacian, see Eq.˜87 and Refs. [46, 47].
VI.1 Fokker–Planck operator
The link between the Langevin and Fokker–Planck pictures is straightforward for an Itō diffusion [51, 52]:
$\mathrm{d}\boldsymbol{X}(t)=\boldsymbol{b}(\boldsymbol{X}(t),t)\,\mathrm{d}t+\sigma(\boldsymbol{X}(t),t)\,\mathrm{d}\boldsymbol{W}(t)\,$, (54)
where $\boldsymbol{X}(t)$ is vector-valued in $\mathbb{R}^{d}$, $\boldsymbol{b}(\cdot,t)$ is a $d$-dimensional, time-dependent vector field, and $\boldsymbol{W}(t)$ is a standard Wiener process in $n$ dimensions scaled by the tensor field $\sigma(\cdot,t)$, which is $d\times n$-matrix-valued and also may (nicely) depend on time. Then, the Fokker–Planck operator corresponding to Eq.˜54 reads
$\mathcal{L}_{\boldsymbol{X}}(t)f(\boldsymbol{x})=-\nabla\cdot[\boldsymbol{b}(\boldsymbol{x},t)f(\boldsymbol{x})]\\ +\frac{1}{2}\nabla\nabla^{\top}:\bigl{[}\sigma(\boldsymbol{x},t)\,\sigma(\boldsymbol{x},t)^{\top}f(\boldsymbol{x})\bigr{]}$ (55)
for suitable real-valued functions $f$; the nabla symbol $\nabla=(\partial_{1},\dots,\partial_{d})$ denotes the vectorial derivative w.r.t. the components of $\boldsymbol{x}$.
The form (54) fits the Itō SDE of rotational Brownian motion, Eq.˜9, upon setting $d=n=3$ and identifying $\boldsymbol{X}(t)=\boldsymbol{u}(t)$ and $(2D_{\text{R}})^{1/2}\,\mathrm{d}\boldsymbol{W}(t)=\delta\boldsymbol{\omega}(t)\,\mathrm{d}t$ [Eqs.˜6, 51a and 51b], see also the text after Eq.˜19. To simplify the notation, we have dropped the superscript $\perp$ at the diffusion and friction coefficients. One reads off that
$b(\boldsymbol{u},t)$ $=-\tau_{\text{R}}^{-1}\boldsymbol{u}-\zeta_{R}^{-1}(\boldsymbol{u}\cdot\boldsymbol{J})\,\boldsymbol{T}_{\text{ext}}(\boldsymbol{u},t)\,$,
$\sigma(\boldsymbol{u})$ $=-(\boldsymbol{u}\cdot\boldsymbol{J})(2D_{\text{R}})^{1/2}\,.$ (56)
It remains to evaluate Eq.˜55 for these choices of $\boldsymbol{b}$ and $\sigma$. After a few calculations to favorably expose the divergence structure of $\mathcal{L}_{\boldsymbol{u}}$ (see Appendix˜B), one finds for the Fokker–Planck operator of rotational Brownian motion, corresponding to Eq.˜9:
$\mathcal{L}_{\boldsymbol{u}}f(\boldsymbol{u})=-\nabla\cdot[\boldsymbol{j}(\boldsymbol{u},t)f(\boldsymbol{u})]\,$, (57)
in terms of the probability flux operator $\boldsymbol{j}(\boldsymbol{u},t)$ separated into its drift and diffusive contributions [cf. Eq.˜82]:
$\boldsymbol{j}(\boldsymbol{u},t)$ $:=\boldsymbol{j}_{\text{drift}}(\boldsymbol{u},t)+\boldsymbol{j}_{\text{diff}}(\boldsymbol{u})\,$,
$\boldsymbol{j}_{\text{drift}}(\boldsymbol{u},t)$ $:=-\zeta_{R}^{-1}(\boldsymbol{u}\cdot\boldsymbol{J})\,\boldsymbol{T}_{\text{ext}}(\boldsymbol{u},t)\,$,
$\boldsymbol{j}_{\text{diff}}(\boldsymbol{u})$ $:=-D_{\text{R}}|\boldsymbol{u}|^{2}\mathcal{Q}_{\boldsymbol{u}}\nabla\,.$ (58)
As above, $\mathcal{Q}_{\boldsymbol{u}}:=\mathds{1}-|\boldsymbol{u}|^{-2}\,\boldsymbol{u}\otimes\boldsymbol{u}$ is the orthogonal projector onto the plane perpendicular to $\boldsymbol{u}$.
Normalization of $\boldsymbol{u}$ must not be assumed at this point since Eq.˜55 is an unconstrained PDE problem in $\mathbb{R}^{3}$; rather it needs to be proven that the probability flux generated by $\mathcal{L}_{\boldsymbol{u}}(t)$ preserves the norm of $\boldsymbol{u}$. Both fluxes are obviously perpendicular to the orientation, $\boldsymbol{u}\cdot\boldsymbol{j}_{\text{drift}}=0$ and $\boldsymbol{u}\cdot\boldsymbol{j}_{\text{diff}}=0$, due to the cross product in the drift part and the projector $\mathcal{Q}_{\boldsymbol{u}}$ in the diffusive part. We conclude that if $|\boldsymbol{u}_{0}|=1$ the solution $p(\boldsymbol{u},t|\boldsymbol{u}_{0},t_{0})$ of Eq.˜53 remains concentrated on the unit sphere under time evolution; this holds also true if one uses a normalized initial distribution rather than a fixed value $\boldsymbol{u}_{0}$. Under this condition, we can put $|\boldsymbol{u}|=1$ in Eqs.˜57 and 58 so that $\mathcal{L}_{\boldsymbol{u}}(t)$ is actually a differential operator that is restricted to the unit sphere. The latter fact is evident in a representation of $\mathcal{L}_{\boldsymbol{u}}(t)$ in terms of spherical coordinates, which contains only derivatives w.r.t. the polar and azimuthal angles $(\vartheta,\varphi)$, but not the norm $\varrho=|\boldsymbol{u}|$ [Eqs.˜87 and 89]. Hence, the norm of $\boldsymbol{u}$ is conserved under the time evolution generated by $\mathcal{L}_{\boldsymbol{u}}(t)$.
VI.2 Detailed balance
We emphasize that, in the absence of the external torque, Eq.˜57 does not contain a drift term, $\boldsymbol{j}_{\text{drift}}=0$, and describes unbiased, free rotational diffusion. The term $\tau_{\text{R}}^{-1}\boldsymbol{u}$ in $\boldsymbol{b}(\boldsymbol{u},t)$ is merely an apparent drift and does not contribute to the probability flux, due to $\tau_{\text{R}}=2D_{\text{R}}$. One verifies that $\mathcal{L}_{\boldsymbol{u}}$ indeed coincides with the expected free diffusion operator on the sphere, i.e., the angular part of the Laplacian multiplied by $D_{\text{R}}$ [Eq.˜87], which is derived in Appendix˜B. It follows that the only stationary distribution, solving $\mathcal{L}_{\boldsymbol{u}}\bar{p}(\boldsymbol{u})=0$, is the uniform distribution, $\bar{p}(\boldsymbol{u})=1/4\pi$, which also satisfies detailed balance, $\boldsymbol{j}_{\text{diff}}(\boldsymbol{u})\,\bar{p}(\boldsymbol{u})=0$, and, hence, describes equilibrium.
More generally, if the external torque arises from a time-independent potential energy $V(\boldsymbol{u})$ such that
$(\boldsymbol{u}\cdot\boldsymbol{J})\boldsymbol{T}_{\text{ext}}(\boldsymbol{u})=|\boldsymbol{u}|^{2}\mathcal{Q}_{\boldsymbol{u}}[\nabla V(\boldsymbol{u})]\,$, (59)
the following holds: The probability fluxes (58) satisfy detailed balance,
$\boldsymbol{j}_{\text{drift}}(\boldsymbol{u})+\boldsymbol{j}_{\text{diff}}(\boldsymbol{u})]\,p_{\text{eq}}(\boldsymbol{u})=0\,$, (60)
if and only if the distribution $p_{\text{eq}}(\boldsymbol{u})$ assumes the Boltzmann form,
$p_{\text{eq}}(\boldsymbol{u})\propto\exp\boldsymbol{(}{-\beta V(\boldsymbol{u})}\boldsymbol{)}\,$, (61)
characteristic of equilibrium at inverse temperature $\beta:=(k_{\text{B}}T)^{-1}=(D_{\text{R}}\zeta_{R})^{-1}$. Such $p_{\text{eq}}(\boldsymbol{u})$ is trivially also a stationary solution, i.e., $\mathcal{L}_{\boldsymbol{u}}p_{\text{eq}}(\boldsymbol{u})=0$.
The statement is proven by, first, noting the equivalence of Eq.˜60 with
$\mathcal{Q}_{\boldsymbol{u}}\bigl{\{}\beta[\nabla V(\boldsymbol{u})]\,p_{\text{eq}}(\boldsymbol{u})+\nabla p_{\text{eq}}(\boldsymbol{u})\bigr{\}}=0\,.$ (62)
It is obvious that Eq.˜61 satisfies this condition. Conversely, searching for solutions to Eq.˜62, we exploit the positivity of the probability density and write $p_{\text{eq}}(\boldsymbol{u})=\exp\boldsymbol{(}{-\Phi(\boldsymbol{u})}\boldsymbol{)}$, which implies that
$\mathcal{Q}_{\boldsymbol{u}}\bigl{[}\beta\nabla V(\boldsymbol{u})-\nabla\Phi(\boldsymbol{u})\bigr{]}=0\,.$ (63)
Therefore, the term in brackets either vanishes or points parallel to the direction of $\boldsymbol{u}$. It follows that $\Phi(\boldsymbol{u})=\beta V(\boldsymbol{u})+\chi(\boldsymbol{u})$ for some field $\chi(\cdot)$ with $\nabla\chi(\boldsymbol{u})\parallel\boldsymbol{u}$. Hence, $\chi(\boldsymbol{u})$ depends only on the magnitude of $\boldsymbol{u}$, but not its direction, and reduces to a ($|\boldsymbol{u}|$-dependent) constant, which is fixed by the normalization of $p_{\text{eq}}(\boldsymbol{u})$. We conclude that Eq.˜61 is the only solution to Eq.˜60.
VI.3 Uniqueness of the stationary solution
Finally, we show that the equilibrium solution (61) is the only stationary solution $\bar{p}(\boldsymbol{u})$ if the torque is derived from a potential energy $V(\boldsymbol{u})$, see Eq.˜59. For any $\bar{p}(\boldsymbol{u})$ with $\mathcal{L}_{\boldsymbol{u}}\bar{p}(\boldsymbol{u})=0$, we define $q(\boldsymbol{u})=\exp\boldsymbol{(}{\beta V(\boldsymbol{u})}\boldsymbol{)}\,\bar{p}(\boldsymbol{u})$. Combining with Eqs.˜57, 58 and 59, it implies
$\mathcal{L}_{\boldsymbol{u}}\bar{p}(\boldsymbol{u})=\nabla\cdot\bigl{[}D_{R}\,\text{e}^{-\beta V(\boldsymbol{u})}|\boldsymbol{u}|^{2}\mathcal{Q}_{\boldsymbol{u}}\nabla q(\boldsymbol{u})\bigr{]}=0\,;$ (64)
also see Eq.˜62 for the expression in brackets. Without loss of generality, we can fix $|\boldsymbol{u}|=1$. Taking the $L^{2}$-inner product of $q(\boldsymbol{u})$ and $\mathcal{L}_{\boldsymbol{u}}\bar{p}(\boldsymbol{u})$ on the unit sphere $S=\{\boldsymbol{u};|\boldsymbol{u}|=1\}$, we have
$\int_{S}q(\boldsymbol{u})\,\nabla\cdot\bigl{[}D_{R}\,\text{e}^{-\beta V(\boldsymbol{u})}\mathcal{Q}_{\boldsymbol{u}}\nabla q(\boldsymbol{u})\bigr{]}\mathrm{d}\boldsymbol{u}=0\,.$ (65)
By the generalized divergence theorem (Stokes’ theorem for manifolds) [90] and observing that the sphere has no boundary, it follows
$\int_{S}\nabla q(\boldsymbol{u})\,\cdot\bigl{[}D_{R}\,\text{e}^{-\beta V(\boldsymbol{u})}\mathcal{Q}_{\boldsymbol{u}}\nabla q(\boldsymbol{u})\bigr{]}\mathrm{d}\boldsymbol{u}=0\,$, (66)
which can be re-arranged, using $D_{\text{R}}>0$ and the projection property $\mathcal{Q}_{\boldsymbol{u}}^{2}=\mathcal{Q}_{\boldsymbol{u}}$, to read
$\int_{S}\text{e}^{-\beta V(\boldsymbol{u})}\bigl{|}\mathcal{Q}_{\boldsymbol{u}}\nabla q(\boldsymbol{u})\bigr{|}^{2}\,\mathrm{d}\boldsymbol{u}=0\,.$ (67)
The integrand is non-negative, so it must vanish almost everywhere; in particular,
$\mathcal{Q}_{\boldsymbol{u}}\nabla q(\boldsymbol{u})=0\quad\text{for all~}\boldsymbol{u}\in S\,$, (68)
since $\nabla q(\boldsymbol{u})$ is continuous. In the same way as above for $\chi(\boldsymbol{u})$, it follows that $\nabla q(\boldsymbol{u})\parallel\boldsymbol{u}$ and that $q(\boldsymbol{u})$ can depend only on the magnitude $|\boldsymbol{u}|$, which is fixed. This is also seen from writing the operator $\mathcal{Q}_{\boldsymbol{u}}\nabla$ in spherical coordinates [Eq.˜85]. Hence, $q(\boldsymbol{u})=\text{const}$, and we conclude that Eq.˜61 is the only stationary probability distribution in case of a potential torque.
VI.4 Example: dipole in a homogeneous field
A typical application is the alignment of a magnetic dipole $\mu\boldsymbol{u}$ (with $|\boldsymbol{u}|=1$) in a spatially homogeneous and constant magnetic field $\boldsymbol{B}$; see e.g. Ref. [93] for an experiment. The potential energy of the dipole is $V(\boldsymbol{u})=-\mu\boldsymbol{B}\cdot\boldsymbol{u}$, which induces the torque $\boldsymbol{T}_{\text{ext}}(\boldsymbol{u})=\mu\boldsymbol{u}\times\boldsymbol{B}$. It results in [cf. Eq.˜59]
$\boldsymbol{u}\times\boldsymbol{T}_{\text{ext}}(\boldsymbol{u})$ $=-\mu\boldsymbol{B}+\boldsymbol{u}(\boldsymbol{u}\cdot\mu\boldsymbol{B})$
$=\mathcal{Q}_{\boldsymbol{u}}\nabla V(\boldsymbol{u})$ (69)
and, thus, the equilibrium distribution [Eq.˜61]
$p_{\text{eq}}(\boldsymbol{u})=\frac{\mu B/k_{\text{B}}T}{4\pi\sinh(\mu B/k_{\text{B}}T)}\,\exp(\mu\boldsymbol{B}\cdot\boldsymbol{u}/k_{\text{B}}T)\,.$ (70)
Appendix˜C provides equivalent expressions of these results using spherical coordinates alongside with detailed calculations.
VII Summary and conclusions
@@
Brownian motion on the unit sphere is a central ingredient to models of axisymmetric colloidal particles, in which it describes the stochastic evolution of the symmetry axis; examples are oblate and prolate ellipsoids, but also active Brownian particles, e.g., Janus spheres, where such an axis is given by the propulsion direction. Here, we have constructed the corresponding overdamped stochastic process as a generalization of Brownian motion in flat (Euclidean) spaces, specifically, through a sequence of infinitesimal random rotations [Eq.˜10]. The latter construction is known as McKean–Gangolli injection [90], which transforms a stochastic process on the Lie algebra — here, $\mathfrak{so}(3)$ — into the generated Lie group — here, the rotation group $SO(3)$. Within Itō calculus, it is then straightforward to derive a stochastic differential (i.e., Langevin) equation of the process [Eq.˜9]. Transforming the equation to its equivalent form in Stratonovich calculus [Eq.˜7] yields the same equation as is obtained heuristically from the underdamped Langevin equation of rotational Brownian motion [Eqs.˜2 and 4]. In particular, the latter form is free from the apparent, noise-induced drift term present in the Itō form. We have shown that, indeed, this additional term does not generate a drift (bias) of the rotational motion; but rather, it is needed to preserve the normalization of the orientation vector, $|\boldsymbol{u}(t)|=1$, under the stochastic dynamics. We have also derived the Fokker–Planck operator $\mathcal{L}_{\boldsymbol{u}}$ corresponding to the overdamped Langevin equation [Eqs.˜57 and 58]. Provided that the external torque originates from a potential energy [Eq.˜59], the form of $\mathcal{L}_{\boldsymbol{u}}$ has allowed us to prove that the constructed process obeys detailed balance [Eq.˜60] and uniquely yields the expected Boltzmann form of the equilibrium distribution [Eq.˜61], which is the only stationary distribution in this case (Section˜VI.3).
@@
In generic numerical integration schemes of the overdamped Langevin equation (in either form), such as the Euler–Maruyama and Milstein schemes, the normalization constraint is satisfied only asymptotically for vanishing time step, $\Delta t\to 0$}, depending on their strong order of convergence. As an alternative, we have proposed a geometric integration scheme [Eq.˜24], which is based on finite random rotations and which satisfies the constraint exactly for any time step $\Delta t$. Combining this approach with Gaussian distributed random rotation angles yields an immediate and simple implementation of this geometric integrator (Table˜1), which we have shown to converge weakly at order 1, i.e., $O(\Delta t)$, and which can also be combined with a deterministic bias, e.g., due to an external torque or to model circle swimmers. Moreover, we have outlined an improved algorithm that exactly generates the (free) propagator of rotational Brownian motion for an arbitrary time step $\Delta t$ if an advanced sampling of the rotation angle is used [see Eq.˜38]. Finally, we have demonstrated numerically and by deriving the corresponding propagators that the proposed Gaussian geometric scheme converges more rapidly than a widely used projection scheme [Eqs. (39)]. The smaller prefactor of the discretization error allows for an integration time step that is an order of magnitude larger (Fig.˜1); in particular, the dynamics of free rotational diffusion is accurately generated also for time steps as large as $D_{\text{R}}\Delta t\approx 0.3$, i.e., typical rotation angles of $\vartheta\approx\pi/4$ in a single step (Fig.˜2). As a consequence, suitable time steps of the geometric scheme are determined by the physical torques and not the constraint on $|\boldsymbol{u}|$. Exact expressions for the orientational correlation functions [Eq.˜31] and for the equilibrium distribution of rotational Brownian motion [Eq.˜70] may serve as references to test actual implementations of the algorithm; an exemplary implementation is provided in the open-source software HAL’s MD package [94] for hardware-accelerated simulation and analysis of many-particle systems.
Acknowledgements.
Financial support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—MATH+: The Berlin Mathematics Research Center (EXC-2046/1)—Project No. 390685689 (Subprojects AA1-18 and EF4-10) and further under Project No. 523950429 is gratefully acknowledged.
@@ Appendix A Statistics of the projected angular velocity
For a derivation of the first two moments of the projected angular velocity $\omega_{\perp}(t)$, given in Eqs.˜51a and 51b, we switch to the integrated noise process,
@@
which in differential form reads $\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=\boldsymbol{\omega}_{\perp}(t)\,\mathrm{d}t$. It is a proper Itō processes and, using Eqs.˜3 and 50, it obeys the following (Itō) SDE in standard form [51, 52]:
$\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t),t)\,\mathrm{d}t\\ +(2D_{R}^{\perp})^{1/2}\mathcal{Q}_{\boldsymbol{u}(t)}\,\mathrm{d}\boldsymbol{W}(t)\,$, (72)
which is driven by a standard three-dimensional Wiener process $\boldsymbol{W}(t)$ normalized such that $\langle|\boldsymbol{W}(t)|^{2}\rangle=3t$.
@@
$=E\biggl{[}\int_{t}^{t+\Delta t}(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t^{\prime})}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t^{\prime}),t^{\prime})\,\mathrm{d}t^{\prime}\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$\simeq(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t),t)\Delta t\,;\quad\Delta t\to 0.$ (73)
@@
Switching to the derivative of $\boldsymbol{\Omega}_{\perp}(t)$ in Eq.˜73 yields Eq.˜51a.
For the covariance function of $\boldsymbol{\Omega}_{\perp}(t)$, we compensate the drift and consider the rescaled process $\mathrm{d}\boldsymbol{W}^{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}(t)}\,\mathrm{d}\boldsymbol{W}(t)$ such that $\int_{0}^{t}\delta\boldsymbol{\omega}_{\perp}(t^{\prime})\,\mathrm{d}t^{\prime}=(2D_{R}^{\perp})^{1/2}\boldsymbol{W}^{\perp}(t)$. Non-overlapping increments of $\boldsymbol{W}^{\perp}(t)$ are uncorrelated, which is seen by choosing $\Delta t^{\prime},\Delta t>0$ and $t^{\prime}>t+\Delta t$ and performing standard manipulations of conditional expectation [52]:
@@
$=E[E[\Delta W^{\perp}_{i}(t^{\prime})|\mathcal{F}(t^{\prime})]\,\Delta W^{\perp}_{j}(t)|\boldsymbol{u}(t)]$ (75)
@@
$=0\,$, (76)
where we have used the so-called tower property of $E[\cdot]$ in the first step, the Markov property of $\boldsymbol{W}^{\perp}(t)$ in the second, and finally, $E[\Delta W^{\perp}_{i}(t^{\prime})|\boldsymbol{u}(t^{\prime})]=E[\Delta W^{\perp}_{i}(t^{\prime})]=0$ due to the independence of $\Delta W^{\perp}_{i}(t^{\prime})$ and $\boldsymbol{u}(t^{\prime})$; the symbol $\mathcal{F}(t)$ denotes the history of $\boldsymbol{u}$ up to time $t$. Choosing the same time interval for both increments, it holds:
@@
$\simeq[\mathcal{Q}_{\boldsymbol{u}(t)}]_{ij}\Delta t\,;\quad\Delta t\to 0\,.$ (77)